89 children were at the party, when 50 % of the girls and 20 % of the boys left, there were 55 children remaining at the party. How many boys were there at the party at first ?
Answer by malglu(63)

at the start
boys (b) + girls (g) = 89
after some left, there are now
1/2 the original amount of girls and 4/5 of the original amount of boys left, and these equal 55 kids altogether.
if you write these out as simultaneous equations
89 = b+g
55 = 1/2g + 4/5 b
then mutiply the second equation by 2
110 = g + 8/5b
89 = g + b
now minus the second one from the first
21 = 3/5b
multiply by 5, then divdide by 3 to get
b=35